OpenStudy (tiffany_rhodes):

Consider the function g(f(x)) where f: A --> B and g: B --> C. I know that if both f and g are bijective, then g(f(x)) is also bijective and therefore has an inverse. But what if only f is bijective? or if only g is bijective? What does that imply about g(f(x)) ?

OpenStudy (zzr0ck3r):

Did I not answer this already?

OpenStudy (zzr0ck3r):

\(f(x)=x\) is a bijection \(g(x)=x^2\) is not a bijection \(f(g(x))=g(f(x))= x^2\) is not a bijection